OpenStudy (anonymous):

hi can you write y= 1/-3 absolute value of x? how do you draw this graph?

OpenStudy (anonymous):

like this: \[y=\frac{1}{-3|x|}\]?

OpenStudy (anonymous):

no y= (1/ -3 ) absolute valute of x

OpenStudy (anonymous):

so \[y=\frac{1}{-3}|x|\] like this?

OpenStudy (anonymous):

yes that's right? can you draw this?

OpenStudy (anonymous):

i write this equation from a graph

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

that's the graph of \[y=\frac{1}{-3}|x|\]

OpenStudy (anonymous):

oh and when draw a positive graph you write how? 

OpenStudy (anonymous):

Then the number in front would not be negative (in this case the 1/-3)

OpenStudy (anonymous):

(h,k)=(0,0) and a=y/x=1/-3? right?

OpenStudy (anonymous):

'V' graph does not have a negative sign outside abs function and before x

OpenStudy (anonymous):

(h,k)=(0,0) and a=y/x=1/-3 are correct

OpenStudy (anonymous):

so my answer is true? with this graph

OpenStudy (anonymous):

yes absolutely true